OpenStudy (blah):

whats 3/m^2 = m-4 / 3m^2 + 2/3m^2

OpenStudy (anonymous):

Common denominator of m^2 and 3m^2 = 3m^2 So multiply every term by 3m^2

OpenStudy (anonymous):

3m^2 * 3/m^2 = 3m^2(m - 4)/3m^2 + 3m^2 * 2/3m^2

OpenStudy (anonymous):

3 * 3 = m - 4 + 2 9 = m - 2 11 = m